OpenStudy (anonymous):

if theaverage weight of 20 women is 130 pounds.and if the average weight of 3/4 of them is 140 pounds. what will b the average weight of remaining 5?? tell me how to solve this question ASAP

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

average is calculated by total weight/total women so (y+z)/20 = 130 3/4 of them indicated 15 women which is given to be 140 pounds so y/15 = 140 the average weight of the remaining 5 would be z/5 = x you can solve for y which is 140/15 so using the first equation: (140/15+z)=130 and the second equation z/5=x -> z=5x then 140/15+5x =130 then you solve for x

OpenStudy (anonymous):

thxxx a lott